OpenStudy (buggiethebug):

Explain in words why the instantaneous rate of change of the function f(x)=|x| cannot be determined at point (0,0)

OpenStudy (anonymous):

because it has a nice corner there

OpenStudy (anonymous):



OpenStudy (anonymous):

rate of change for \(x<0\) is always \(-1\) and for \(x>0\) it is always \(1\)

OpenStudy (buggiethebug):

but it is asking for the instantaneous rate of change?, is it because it's a cusp

OpenStudy (anonymous):

because on one side it is minus one (always) and on the other side it is one

OpenStudy (anonymous):

i guess you can say it is a "cusp" instantaneous rate of change is the derivative right?

OpenStudy (anonymous):

absolute value is a piecewise function \[ = |x| = \left\{\begin{array}{rcc} -x & \text{if} & x \leq 0 \\ x& \text{if} & x >0 \end{array} \right. \]

OpenStudy (buggiethebug):

isnt it still approaching 0 from both sides.. doesnt that mean that there is a limit

OpenStudy (buggiethebug):

you are awesome thank youuu :3